OpenStudy (anonymous):

Two brothers are digging a drainage ditch for a french drain around their home. The ditch is to be 2 feet deep, 1 foot wide, and 40 feet long. The older brother can dig the ditch in 14 hours. The younger brother can dig the ditch in 17 hours. How long will the two brothers working together take to dig the ditch? (Round your answer to the nearest tenth.) @AwesomeB @calculusxy

OpenStudy (calculusxy):

Wait. I am going to write the equation.

OpenStudy (calculusxy):

x=time when working together x/17+x/14=1 14x+17x=31x 31x=238 238/31= 7hours, 40 minutes, 38.7 seconds

OpenStudy (calculusxy):

@shortie18

OpenStudy (anonymous):

the answer is ganna be 7.7 hours

OpenStudy (calculusxy):

Yes

OpenStudy (anonymous):

ok thanks

OpenStudy (calculusxy):

No problem. :)

OpenStudy (anonymous):

i am ganna post another one